OpenStudy (anonymous):

which of the following are polynomals A.0 B.1/x+x C.-xyz^4 D.2x-x^-1 E.x^y select all correct answers

OpenStudy (neer2890):

In mathematics, a polynomial is an expression consisting of variables and coefficients, that involves only the operations of addition, subtraction, multiplication, and non-negative integer exponents.

OpenStudy (anonymous):

so which ones would work in the question

OpenStudy (neer2890):

Polynomials of small degree have been given specific names. A polynomial of degree zero is a constant polynomial or simply a constant. Polynomials of degree one, two or three are respectively linear polynomials, quadratic polynomials and cubic polynomials. The polynomial 0, which may be considered to have no terms at all, is called the zero polynomial.

OpenStudy (anonymous):

so it would be A,B,E

OpenStudy (neer2890):

for e.g., 3xy^-2 is not, because the exponent is "-2" (exponents can only be 0,1,2,...) 2/(x+2) is not, because dividing by a variable is not allowed 1/x is not either √x is not, because the exponent is "½" (see fractional exponents) But these are allowed: x/2 is allowed, because you can divide by a constant also 3x/8 for the same reason √2 is allowed, because it is a constant (= 1.4142...etc)

OpenStudy (anonymous):

so whats the answers to the question

OpenStudy (neer2890):

what do you think are the correct options?

OpenStudy (anonymous):

B,E

OpenStudy (anonymous):

what do u think ?

OpenStudy (neer2890):

why not A. As i mentioned earlier that 0 itself is a zero polynomial.

OpenStudy (anonymous):

so it would just be A

OpenStudy (neer2890):

and why not D 2x -x^-1 2x is a polynomial and x^-1 is not. But still it's partially a polynomial.

OpenStudy (anonymous):

so A,D thats all ?

OpenStudy (anonymous):

wouldnt B work

OpenStudy (neer2890):

No.. the question asks about selecting all the variables, it means that it may contain more than one polynomial.

OpenStudy (neer2890):

yeah.. B is a polynomial.

OpenStudy (anonymous):

So the answers would be A.B,E and cant use D it has to be a polynomial cant just be half

OpenStudy (neer2890):

wait..,why can't C?

OpenStudy (neer2890):

a polynomial is an expression consisting of variables and coefficients and -xyz^4 is a polynomial by definition.

OpenStudy (anonymous):

let f(x)=1 x x^3-3 square root2/2x^99 what is f(0)

OpenStudy (neer2890):

just put 0 inplace of x in f(x) and evalutate it.

OpenStudy (anonymous):

so its 0

OpenStudy (neer2890):

is this your question..? \[f(x)=1*x ^{3}-\frac{ 3\sqrt{2} }{ 2x ^{99} }\]

OpenStudy (anonymous):

ok next ?

OpenStudy (anonymous):

let f(x)=x^3y^12+21.6x^13-sq root 8 A.the function f is a 13th degree trinomial B.the function f is a 12th degree trinomial C.the function f is a 15th degree trinomial D.the function f is a 13th degree binomial E. the function f is not a valid polynomial

OpenStudy (neer2890):

if this is your question, then remember anything divide by zero is indeterminate. so\[1*0^{3}=0\] but \[\frac{ 3\sqrt{2} }{ 0^{99} }\] is indeterminate. So, we can not find f(0). But you can solve this by using limits to find the nearest value.

OpenStudy (neer2890):

To find the degree of polynomial,just add the powers of polynomial terms and whichever is highest is the degree. for e.g., in first term x^3y^12 (x has power 3 and y has 12) so, overall degree of this term is (3+12=15) similarly find the degree of other terms and the highest one is the degree of whole polynomial.

OpenStudy (anonymous):

so the degree is 13 and its trinomial ?

OpenStudy (neer2890):

i think 15 is higher than 13.

OpenStudy (anonymous):

where did u get 15

OpenStudy (neer2890):

i already explained buddy.., try to read that.

OpenStudy (anonymous):

let f(x)=2x^3-7x^3+5x^3+123 A.the function f is a 3rd degree trinomial B.the function f is a 12th degree monomial C.the function f is a 3rd degree bionomial D.the function f is a 0th degree monomial E.the function is not a valid polynomial

OpenStudy (neer2890):

try same procedure as earlier. f(x)=2x^3-7x^3+5x^3+123 =x^3(2-7+5)+123 =x^3(0)+123

OpenStudy (neer2890):

Can you answer it now..?

OpenStudy (anonymous):

i need help ive got 4 more questions can u help me with them i dont have any help at home

OpenStudy (anonymous):

i got 123

OpenStudy (neer2890):

0th degree polynomials are another word for constants. and we know that 123 is a constant and is a monomial also.

OpenStudy (anonymous):

so its D

OpenStudy (neer2890):

yeah,,

OpenStudy (anonymous):

let f(x)=3x^2=x^3-2 A.the leading term is x^3,the constant term is -2 B.the leading term of f is 3x^2,and its contant term is 2 C.the leading term of f is 3x^2,and its constant term is -2 D.the leading term of f is x^3,and its constant term is 2

OpenStudy (anonymous):

i think its C.

OpenStudy (neer2890):

f(x)=3x^2+x^3-2..?

OpenStudy (anonymous):

yes sorry i put = i meant +

OpenStudy (neer2890):

always higher deree terms are arranged on left hand side. means we start a polynomial from left to right in decreasing degree. so f(x)=x^3+3x^2-2 that means the leading term is x^3 and constant is -2.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the polynomial functions f(x)=12x^2+5x^4-2and g(x)-2+12x^2+5x^4 are equal A. true B.falsee

OpenStudy (anonymous):

the polynomial functions f(x)=12x^2+5x^4-2and g(x)-2+12x^2+5x^4 are equal A. true B.false

OpenStudy (anonymous):

be back in 15

OpenStudy (neer2890):

i already told you buddy..., polynomials are always arranged from left to right in order of descending degree of terms. so x^3 is leftwards than x^2 and x and constants. x^4 remains leftwards than x^3 and all powers of x less than 3

OpenStudy (neer2890):

so f(x)=12x^2+5x^4-2 =5x^4+12x^2-2 and g(x)=-2+12x^2+5x^4 =5x^4+12x^2-2 so f(x)=g(x)

OpenStudy (anonymous):

so is true

OpenStudy (neer2890):

yeah

OpenStudy (anonymous):

can u help me with factors

OpenStudy (anonymous):

what value in place of the question mark ,makes the polynomial factor a perfect square x^2+12x+?